# SReferences

### v1

@@ IV-D1 Influence of Different Components
We show in Tab. III that hyperbolic spaces are better calibrated and a mapping of the final layer to this space increases the performance after base training by $0.79$pp which complements findings in [weber2024flattening]. Further, the performance improvement increases to $2.52$pp in mIoU after the final step which can be attributed to a difference of $5.73$pp in novel classes. Consequently, we suggest that equidistant class mappings in hyperbolic space support incremental learning. Further, we find that our hierarchical loss results in a performance improvement of $1.85$pp which can be attributed to a performance increase in base and novel classes. Base classes benefit from the rigidity of the feature space as well as a high quality of learned features in base training ($1.34$pp in comparison to the flat classifier) which helps prevent forgetting [Mittal_2021_CVPR]. Additionally, the generalization on new classes is amplified by $6.52$pp. We reason that learning new classes is eased in a hierarchy as their mapping is already defined by trained ancestor nodes. Consequently, we motivate hierarchical modeling in neural networks for open-world learning. Our regularization losses increase the performance by $0.21$pp and $0.14$pp, respectively. We note that newly added classes significantly benefit from our constant sparsity constraint ($0.73$pp) as their relative frequency largely differs in subsequent incremental steps. Further, constraining implicit class relations further improves the knowledge retention of new classes by $0.41$pp.
Table III: Ablation study on the efficacy of various components of TOPICS . All results are reported on Cityscapes in mIoU (%). The setting $1-14_{0}$ represents the mIoU on base classes at time $t=0$.
@@
Space $\mathcal{L}_{hier}$ $\mathcal{L}_{dist}$ $\mathcal{L}_{rel}$ 1-140 1-14 15-19 all
Eucl. 72.72 71.06 29.09 57.02
@@ IV-D2 Learning Rate and Epochs
Some benchmarks adopt a lower learning rate [Douillard2020PLOPLW, cermelli2020mib, baek2022_dkd, goswami2023attribution] and train for less epochs [Douillard2020PLOPLW, goswami2023attribution, cermelli2020mib] during incremental steps. While our model benefits from the selected settings, we additionally show in Tab. IV that TOPICS also outperforms all benchmarks with their selected hyperparameters. We note that training our model for shorter epochs or a lower learning rate significantly decreases the generalization on novel classes ($\mathcal{C}_{1:T}$). Further, the performance on base classes is affected which we reason with false positive segmentation of base classes. When training for $30$ epochs on a learning rate of $0.01$ this effect seems diminished as the base performance peaks with this reduced number of iterations.
@@
In this supplementary material, we present extended results on Cityscapes and Mapillary Vistas 2.0 in Sec. S.1. Further, we show additional illustrations of the applied taxonomic trees in Sec. S.2.
@@
We perform a detailed analysis of the base IoU scores shown in Sec. IV-C. All compared methods achieve different performances after base training which significantly influences the observed performance on base classes after the final incremental step. Consequently, we note the relative performance degradation of base classes as a percentage of their initial performance in Tab. S.1 and Tab. S.2. On Cityscapes, we observe that TOPICS presents the least relative performance drop on incremental settings from the background. For the incremental setting from known classes, baseline methods such as DKD [baek2022_dkd] and MicroSeg[zhang22_microseg] can similarly retain knowledge on the non-incremented base class ’sky’ which we reason with their missing plasticity to generalize to novel classes.
@@
S.2 Extended Quantitative Results
@@
Figure S.0: Visualization of the Cityscapes class taxonomic tree $\mathcal{H}_{6}$. Base classes ($\mathcal{C}^{1}$) are colored in black (ancestors) and blue (leaf classes) whereas novel classes ($\mathcal{C}^{2:T}$) are colored in red.
@@
Figure S.1: Visualization of the Cityscapes class taxonomic tree $\mathcal{H}_{4}$. Base classes ($\mathcal{C}^{1}$) are colored in black (ancestors) and blue (leaf classes) whereas novel classes ($\mathcal{C}^{2:T}$) are colored in red for Cityscapes 15-1(6 tasks).
@@
Figure S.2: Visualization of the Cityscapes class taxonomic tree $\mathcal{H}_{3}$. Base classes ($\mathcal{C}^{1}$) are colored in black (ancestors) and blue (leaf classes) whereas novel classes ($\mathcal{C}^{2:T}$) are colored in red for Cityscapes 15-1(6 tasks).
@@
Figure S.3: Visualization of the Mappillary Vistas class taxonomic tree $\mathcal{H}_{6}$. Base classes ($\mathcal{C}^{1}$) are colored in black (ancestors) and blue (leaf classes) whereas novel classes ($\mathcal{C}^{2:T}$) are colored in red for Mapillary 71-10(5 tasks). The tree is subdivided at the root for visualization purposes.

### v2

@@ IV-D1 Influence of Different Components
We show in Tab. III that pseudo-labeling of the background prevents the collapse of base class performance when training with sigmoid activations and binary cross-entropy on flat labels. Further, hyperbolic spaces are better calibrated as the performance after base training increases by $0.79$pp which complements findings in [weber2024flattening]. The performance improvement increases to $2.52$pp in mIoU after the final step which can be attributed to a difference of $5.73$pp in novel classes. Consequently, we suggest that equidistant class mappings in hyperbolic space support incremental learning. Further, we find that our hierarchical loss results in a performance improvement of $1.85$pp which can be attributed to a performance increase in base and novel classes. Base classes benefit from the rigidity of the feature space as well as a high quality of learned features in base training ($1.34$pp in comparison to the flat classifier) which helps prevent forgetting. Additionally, the generalization on new classes is amplified by $6.52$pp. We reason that learning new classes is eased in a hierarchy as their mapping is already defined by trained ancestor nodes. Thus, we motivate hierarchical modeling in neural networks for open-world learning. Our regularization losses increase the performance by $0.21$pp and $0.14$pp, respectively. Newly added classes significantly benefit from our constant sparsity constraint ($0.73$pp) as their relative frequency largely differs in incremental steps. Constraining implicit class relations further improve the retention of new classes by $0.41$pp. Further results are in the supplementary material Sec. S.5.
Table III: Ablation study on the efficacy of various components of TOPICS . All results are reported on Cityscapes in mIoU (%). The setting $1-14_{0}$ represents the mIoU on base classes at time $t=0$. PL: pseudo-labeling of the background with the old model.
@@
Space $PL$ $\mathcal{L}_{hier}$ $\mathcal{L}_{dist}$ $\mathcal{L}_{rel}$ 1-140 1-14 15-19 all
Eucl. 72.72 0.00 14.55 3.64
72.72 71.06 29.09 57.02
@@ IV-D2 Learning Rate and Epochs
Some benchmarks adopt a lower learning rate [Douillard2020PLOPLW, cermelli2020mib, baek2022_dkd, goswami2023attribution] and train for less epochs [Douillard2020PLOPLW, goswami2023attribution, cermelli2020mib] during incremental steps to prevent forgetting. While our model benefits from the selected settings, we additionally show in Tab. IV that TOPICS also outperforms all benchmarks with their selected hyperparameters. We note that training our model for shorter epochs or a lower learning rate significantly decreases the generalization on novel classes ($\mathcal{C}_{1:T}$). Further, the performance on base classes is affected which we reason with false positive segmentation of base classes. When training for $30$ epochs on a learning rate of $0.01$ this effect seems diminished as the base performance peaks with this reduced number of iterations.
@@
Table S.1: Influence of different class taxonomy trees ($\mathcal{H}$) for results on Cityscapes in mIoU (%). The setting $1-14_{0}$ represents the mIoU on base classes at time $t=0$. The relative decrease in performance on $\mathcal{C}^{1}$ is recorded in (%).
14-1 (6 tasks)
$\mathcal{H}$ 1-140 $\downarrow$1-140 [%] 1-14 15-19 all
$\mathcal{H}_{3}a$ 69.05 4.10 66.22 29.17 56.95
$\mathcal{H}_{4}a$ 69.61 2.99 67.53 34.30 59.22
$\mathcal{H}_{5}a$ 69.35 3.03 67.25 30.59 58.08
$\mathcal{H}_{6}$ 74.84 2.42 73.03 42.47 61.74
Table S.2: Influence of different class taxonomy trees ($\mathcal{H}$) for computational cost on Cityscapes in mIoU (%).
14-1 (6 tasks)
$\mathcal{H}$ training time on GPU average inference time in seconds
$\mathcal{H}_{2}$ 4 hours 43 minutes 0.22
$\mathcal{H}_{4}$ 6 hours 49 minutes 0.22
$\mathcal{H}_{6}$ 8 hours 16 minutes 0.24
@@
Table S.5: Influence of different class taxonomy trees ($\mathcal{H}$) for performance on incremental classes over training increments on Cityscapes in mIoU (%).
14-1 (6 tasks)
$\mathcal{H}_{4}$ $\mathcal{H}_{6}$
Task 15 16 17 18 19 all 15 16 17 18 19 all
2 48.70 - - - - 67.93 61.89 - - - - 69.10
3 34.61 51.43 - - - 65.83 58.69 66.11 - - - 68.16
4 15.04 8.25 14.46 - - 59.25 31.67 31.48 18.12 - - 61.84
5 6.26 3.47 14.39 59.42 - 58.05 27.31 31.17 18.17 63.13 - 61.55
6 7.87 3.92 14.93 53.47 71.99 58.37 28.11 34.43 19.93 56.70 73.15 61.74
In this supplementary material, we present extended results on Cityscapes and Mapillary Vistas 2.0 in Sec. S.1. Further, we show additional illustrations of the applied taxonomic trees in Sec. S.2. Sec. S.3 and Sec. S.4 provide results on additional class taxonomies on Cityscapes and give indications of their associated computational costs. We emphasize the positive impact of our regularization losses with additional experiments on Cityscapes and PascalVOC in Sec. S.5. Last, we show the impact of adding our hierarchical loss function to other baseline methods in Sec. S.7 and further analyze the impact of the margin loss in Sec. S.6.
@@
We perform a detailed analysis of the base IoU scores shown in Sec. IV-C. All compared methods achieve different performances after base training which significantly influences the observed performance on base classes after the final incremental step. Thus, we note the relative performance degradation of base classes as a percentage of their initial performance in Tab. S.3 and Tab. S.4. On Cityscapes, we observe that TOPICS presents the least relative performance drop on incremental settings from the background. For the incremental setting from known classes, DKD [baek2022_dkd] and MicroSeg[zhang22_microseg] can similarly retain knowledge on the non-incremented base class ’sky’ which we reason with their missing plasticity to generalize to novel classes.
@@
S.2 Illustrations of the applied taxonomic trees
@@
Figure S.0: Visualization of the Cityscapes class taxonomic tree $\mathcal{H}_{6}$. Base classes ($\mathcal{C}^{1}$) are colored in black (ancestors) and blue (leaf classes) whereas novel classes ($\mathcal{C}^{2:T}$) are colored in red for Cityscapes 14-1 (6 tasks).
@@
Figure S.1: Visualization of the Cityscapes class taxonomic tree $\mathcal{H}_{4}$. Base classes ($\mathcal{C}^{1}$) are colored in black (ancestors) and blue (leaf classes) whereas novel classes ($\mathcal{C}^{2:T}$) are colored in red for Cityscapes 14-1 (6 tasks).
@@
Figure S.2: Visualization of the Cityscapes class taxonomic tree $\mathcal{H}_{2}$. Base classes ($\mathcal{C}^{1}$) are colored in black (ancestors) and blue (leaf classes) whereas novel classes ($\mathcal{C}^{2:T}$) are colored in red for Cityscapes 14-1 (6 tasks).
Figure S.3: Visualization of the Mappillary Vistas class taxonomic tree. Base classes ($\mathcal{C}^{1}$) are colored in black (ancestors) and blue (leaf classes) whereas novel classes ($\mathcal{C}^{2:T}$) are colored in red for Mapillary 71-10 (5 tasks). The tree is subdivided at the root for visualization purposes.
S.3 Extended Ablation on Influence of different class taxonomy trees
We complement our findings in the paper that deeper hierarchies prevent forgetting with three further taxonomies shown in Fig. S.4, Fig. S.5 and Fig. S.6. For $\mathcal{H}_{3a}$, all predicted classes are at the same hierarchy level (level $3$) in the taxonomy tree. In comparison $\mathcal{H}_{4a}$, extends $\mathcal{H}_{3}a$ but restricts the hierarchy to have a maximum of $3$ sibling nodes. $\mathcal{H}_{5a}$ allows a maximum of $2$ sibling nodes.
We show in Tab. S.1 that $\mathcal{H}_{4a}$ outperforms $\mathcal{H}_{3a}$ and $\mathcal{H}_{5a}$. Consequently, we find that having all relevant classes on the same hierarchy level is not required which proves the suitability of our method for incremental learning with unknown taxonomies. Further, we find that class taxonomies that vary the number of siblings between $2-3$, such as $\mathcal{H}_{4a}$ and our best-performing $\mathcal{H}_{6}$ (Fig. 6), should be preferred. In summary, we propose that the depth of the hierarchy should be correlated to the number of classes in a way that every internal node has $2-3$ child classes and motivate future research to further explore methods that can predict the best-fitted taxonomic hierarchies.
We further analyze the performance differences between $\mathcal{H}_{4}$ and $\mathcal{H}_{6}$ as shown in Tab. LABEL:tab:tree. While the closed-set performance only differs by $0.52$pp, the performance on novel classes is significantly boosted ($12.03$pp) when leveraging $\mathcal{H}_{6}$. We present the final performance on each individual novel class in Tab. S.5. We note that forgetting not only occurs for base classes but also imposes a major challenge for new classes, which are learned at earlier increments. Our proposed hierarchy $\mathcal{H}_{6}$ significantly aids in achieving a higher performance on novel classes but also preserves this performance. Thus, we conclude that deeper hierarchies prevent forgetting in incremental learning.
S.4 Extended Ablation on Computational Costs of different hierarchies
We measure the computational cost as the total incremental training time on two NVIDIA A40’s and average inference time per sample. We present in Tab. S.2 that deeper hierarchies lead to higher training times while the inference time only has a minor increase. This observation can be attributed to the increased complexity of the hierarchical loss with deeper hierarchies.
S.5 Extended Ablation on Impact of our Regularization Losses
We highlight that pseudo-labeling of the background and our taxonomy-aware loss provide very strong regularization for autonomous driving datasets where certain classes, e.g. sky and road, are commonly observed in every image. Consequently, we present results of TOPICS when only pseudo-labeling the background with a probability of 10% during incremental training in Tab. S.6. We disregard the areas of neglected pseudo-labels in all loss function to ensure consistent labeling. We find that the mIoU performance on novel classes drops significantly while their accuracy scores remain high which indicates that novel classes are commonly predicted for ignored image regions. Nevertheless, we find that our regularization losses achieve a performance improvement of $1.37$pp in this experimental setting which can be attributed to an improvement in base and novel classes. Consequently, we highlight that our novel loss functions effectively prevent catastrophic forgetting, which is more pronounced when prior objects appear less frequently in the background.
To strengthen our point, we additionally present results of TOPICS on the PascalVOC dataset for which prior classes don’t commonly appear in other incremental steps in Tab. S.7. We find that the performance improvement of applying our hierarchical loss amounts $6.49$pp in comparison to a binary cross-entropy on flat labels even if we did not tune the hierarchical loss to achieve a comparable performance after base training. The performance improves by additional $28.78$pp with the application of our two regularization losses. This improvement can be attributed to enhanced scores for both base and novel classes.
Table S.6: Ablation study on the efficacy of various components of TOPICS. All results are reported on Cityscapes in mIoU (%). All variants apply pseudo-labeling of the background with the old model with a probability of 10%.
14-1 (6 tasks)
$\mathcal{L}_{hier}$ $\mathcal{L}_{dist}$ $\mathcal{L}_{rel}$ 1-14 15-19 all
62.90 6.63 45.69
64.44 7.08 46.88
62.68 6.63 45.53
64.69 7.13 47.06
Table S.7: Ablation study on the efficacy of various components of TOPICS. All results are reported on PascalVOC in mIoU (%). The setting $1-15_{0}$ represents the mIoU on base classes at time $t=0$. All variants apply pseudo-labeling of the background with the old model.
15-1 (6 tasks)
$\mathcal{L}_{hier}$ $\mathcal{L}_{dist}$ $\mathcal{L}_{rel}$ 1-150 1-15 16-20 all
79.78 18.21 14.15 20.30
75.68 32.31 13.72 26.79
75.68 35.95 13.45 31.67
75.68 67.44 16.03 52.58
75.68 68.59 20.66 55.57
S.6 Extended Ablation on Impact of Margin Loss
The margin-based loss in [liulei24] uses a threshold to determine the minimum distance in its triplet loss function. As suggested by the authors, we tune this parameter according to the depth of the hierarchy. We present further results on the margin loss in Tab. S.8. While the margin loss shows superior results on closed-set training in Euclidean space (column 1-140), this loss function does not prevent forgetting in incremental learning. Further, we do not record any performance benefits of applying this loss to shallow or deep hierarchies in hyperbolic space. We reason this observation with non-linear increasing distances from the center of the Poincaré ball, whereas one fixed distance threshold is not sufficient to promote further alignment of the hierarchy. This finding provides insights into the construction of hyperbolic loss functions.
S.7 Extended Ablation on Impact of our Hierarchical Loss
Further, we provide additional experiments in Tab. S.9 where we add the tree-min loss function to two presented baselines. While the performance of MiB [cermelli2020mib] is boosted by $1.04$pp, the performance on DKD [baek2022_dkd] decreased by $5.34$pp. Consequently, we find that naively integrating taxonomies into existing class-incremental learning methods does not automatically result in improved performance. It requires proper design and alignment with knowledge retention methods to make effective use of this supervision.
Table S.8: Influence of margin loss functions for results on Cityscapes in mIoU (%). The setting $1-14_{0}$ represents the mIoU on base classes at time $t=0$.
14-1 (6 tasks)
Space $\mathcal{H}$ Tree-min Margin Loss 1-140 1-14 15-19 all
Euclidean $\mathcal{H}_{2}$ 68.37 59.37 24.93 47.79
69.69 58.85 25.31 47.52
Hyperb. $\mathcal{H}_{2}$ 66.46 60.86 32.57 50.75
66.31 60.43 30.13 49.83
$\mathcal{H}_{6}$ 74.85 73.03 42.47 61.74
74.20 72.97 35.31 59.91
Table S.9: Integration of hierarchical supervision to baseline methods on Cityscapes in mIoU (%).
14-1 (6 tasks)
Method 1-14 15-19 all
DKD [baek2022_dkd] 68.83 14.70 51.86
DKDhier 61.01 15.27 46.52
MiB [cermelli2020mib] 66.37 14.36 50.05
MiBhier 68.29 13.14 51.09
Figure S.4: Visualization of the Cityscapes class taxonomic tree $\mathcal{H}_{3a}$. Base classes ($\mathcal{C}^{1}$) are colored in black (ancestors) and blue (leaf classes) whereas novel classes ($\mathcal{C}^{2:T}$) are colored in red for Cityscapes 14-1 (6 tasks).
Figure S.5: Visualization of the Cityscapes class taxonomic tree $\mathcal{H}_{4a}$. Base classes ($\mathcal{C}^{1}$) are colored in black (ancestors) and blue (leaf classes) whereas novel classes ($\mathcal{C}^{2:T}$) are colored in red for Cityscapes 14-1 (6 tasks).
Figure S.6: Visualization of the Cityscapes class taxonomic tree $\mathcal{H}_{5a}$. Base classes ($\mathcal{C}^{1}$) are colored in black (ancestors) and blue (leaf classes) whereas novel classes ($\mathcal{C}^{2:T}$) are colored in red for Cityscapes 14-1 (6 tasks).